# Investigating the Representation of Backchannels and Fillers in Fine-tuned Language Models

### v1

@@ D.2 Analysis of the Generation of Backchannels and Fillers
D.3 Extra t-SNE Visualisations
The additional t-SNE figures are shown in Figure 7 to 16. One thing we noticed is that in some of the t-SNE results from LLaMA-3 8B, even in the absence of fine-tuning and contextual information, the distinction between different backchannels/fillers is quite clear. We believe this is the case because the LLaMA models have already learnt some information about backchannels/fillers during their pre-training phase. Another important issue is that, even with fine-tuning, the use of contextual information introduces uncertainty, as some of the visualizations with fine-tuning settings may still have border between different backchannels/fillers.
D.4 Silhouette Scores Statistics and K-Mean Values Before and After PCA Dimensionality Reduction
Tables  6 to 12 respectively show the silhouette scores of each backchannels/fillers in both the original dimensional space and after dimensionality reduction to 100 dimensions using Principal Component Analysis (PCA), along with the average silhouette scores for English and Japanese data. A general trend is that, with fine-tuning for most of the case silhouette score increases. When there is no fine-tuning, adding context size will also have the same effect. However, combining context size with fine-tuning as well as increasing context size will lower the silhouette score. Figure 25 and Table 14 to 21 summarize the change of the k value before and after fine-tuning.
Dialogue 1 Input: /A <ds>yeah</ds> /A /B so we went in and the financed the car for five and we didn’t know that we got kind of taken just a little bit so we kind of upset that we having to you know pay this fifth year on the car because it was just not a wise thing but we learned a good lesson so you know /B Ground truth: /A <ds>well</ds> five years is not that <ds>uh</ds> five years is not that bad of a problem only because you paying more interest but your payments are lower /A Output (no_ft_one): /A this all the experience that we got in that situation yeah I just want to say that it was not just the car that we had to deal with but it was /A Output (ft_one): /A <ds>yeah</ds> we were just <ds>um</ds> my my parents bought the car that we would drive in high school and they put like two hundred down and and I… /A
(a) Example 1
Dialogue 2 Input: /A but then <ds>uh</ds> we also didn’t do spend <ds>uh</ds> time with them trying to trying to <ds>uh</ds> you know just family time /A /B <ds>uh-huh</ds>/B Ground truth: /A <ds>uh</ds> we had a at a fairly young age we started teaching them to play games and card games and so forth and and <ds>uh</ds> <ds>uh</ds> but we didn’t <ds>uh</ds> do it as much or as often as <ds>uh</ds> I would have liked to I would like to do that as a as a family and /A Output (no_ft_one): /A with the family that we have now with the family that we have now. What are you going to do with that kind of logic? I mean, I’d like to solve this problem. But I can’t. Can you help me? I have a little bit of time before the family comes over. So, can you give… /A Output (ft_one): /A I think <ds>uh</ds> my father was really into them and we would you know go camping and stuff and <ds>uh</ds> you know I think if you do that then it important to spend some time with them you know /A /B <ds>uh-huh</ds> /B
(b) Example 2
Dialogue 3 Input: /A so it becomes <ds>uh</ds> a general battle of the w- sort of the election of the lesser of two evils I think so /A /B <ds>yeah</ds> /B Ground truth: /A I don’t know I just <ds>uh</ds> I unhappy with it but and well I I sort of semi-fortunate right now I a graduate student so I don’t make that much /A Output (no_ft_one): /A The concept of "lesser of two evils" suggests that when faced with a difficult decision, it’s better to choose the option that is less undesirable. In the context of the 202 /A Output (ft_one): /A it always seems to be that way <ds>so</ds> <ds>um</ds> what do you think is the other major issue now /A /B <ds>um</ds> I I don’t know I think…/B
(c) Example 3
Dialogue 4 Input: /A <ds>yeah</ds> <ds>well</ds> my <ds>uh</ds> my <ds>uh</ds> probably one of the biggest decisions I think that was very strengthening for our family was rather than have one child make that decision /A /B <ds>uh-huh</ds> /B Ground truth: /A than just delegate it I think that they they had a great deal of <ds>um</ds> all the brothers and sisters got together and they actually had a conference /A Output (no_ft_one): /A one year. For one of one or a rather long time I am one I do one I did one I have one I will one one I had one /A Output (ft_one): /A I think we went to church and I think we sought <ds>um</ds> the advice of those that could see it from a different perspective /A
(d) Example 4
Figure 6: Illustrative examples of generated backchannels/fillers from LLaMA-3.
Table 4: Evaluation metrics for generated backchannels/fillers in English NLG task
@@ D.4 Silhouette Scores Statistics and K-Mean Values Before and After PCA Dimensionality Reduction
Perplexity $\Downarrow$ 197.67 5.30 202.06 91.32 158.64 6.98
BERTScore (F1 %) $\uparrow$ 78.69% 79.99% 76.02% 79.73% 79.67% 79.99%
BLEUScore $\uparrow$ 0.0600 0.0697 0.0698 0.0800 0.0544 0.0731
Table 5: Evaluation metrics for generated backchannels/fillers in Japanese NLG task
@@
Perplexity $\Downarrow$ 255.94 28.51 748.63 90.17 195.36 53.51
BERTScore (F1 %) $\uparrow$ 62.68% 66.31% 63.55% 63.39% 62.96% 63.75%
BLEUScore $\uparrow$ 0.00005 0.00030 0.00021 0.00036 0.00019 0.00026
D.5 Confusion Matrices for the Top 15 Backchannels and Fillers in English and Japanese
As our further prove to the result that fine-tuning helps LMs to learn the representations of backchannels/fillers, we report the distance among different backchannels/fillers before and after fine-tuning in both English and Japanese using confusion matrices (Figures 17 to 24, darker colour indicates larger distance). The fine tuning strategy is NTP and MASK. For this additional experiment, we first obtain the centroids of the k-means clustering corresponding to the maximum silhouette score. Each centroid is an embedding value. After getting the centroids for each backchannels/fillers, we calculate the average of the centroids to get an embedding value which can represent the corresponding backchannels/fillers. We calculate the distance between two backchannels/fillers using euclidean distance. In the final visualisation, we can observe that, in GPT-2 based models, some backchannels/fillers are getting closer after fine-tuning (e.g., ‘of course’ and ‘okay’ in English ) while some are more away from each other (e.g., ‘mmhmm’ and ‘yes’, ‘no’)(Figures 19(a) and 19(b)). In the LLaMA-3 8B based models, some of the backchannels/fillers are considered more similar to each other before fine-tuning and the situation has changed after fine-tuning (Figures 17 and 21). And the same is true for Qwen-3 (Figures 18 and 22) and BERT based models (Figures 20 and 24).
@@ D.5 Confusion Matrices for the Top 15 Backchannels and Fillers in English and Japanese
Figure 7: The t-SNE plots of the backchannels/fillers embeddings from LLaMA-3 model (NTP) under no context setting.
@@
Figure 8: The t-SNE plots of the backchannels/fillers embeddings from LLaMA-3 model (NTP) under full context setting.
@@
Figure 9: The t-SNE plots of the backchannels/fillers embeddings from Qwen-3 model (NTP) under no context setting.
@@
Figure 10: The t-SNE plots of the backchannels/fillers embeddings from Qwen-3 model (NTP) under one context setting.
@@
Figure 11: The t-SNE plots of the backchannels/fillers embeddings from Qwen-3 model (NTP) under full context setting.
@@
Figure 12: The t-SNE plots of the backchannels/fillers embeddings from GPT-2 model (NTP) underthe no context setting.
@@
Figure 13: The t-SNE plots of the backchannels/fillers embeddings from GPT-2 model (NTP) under one context setting.
@@
Figure 14: The t-SNE plots of the backchannels/fillers embeddings from GPT-2 model (TTP) under no and one context setting.
@@
Figure 15: The t-SNE plots of the backchannels/fillers embeddings from BERT model (MASK) under no context setting.
@@
Figure 16: The t-SNE plots of the backchannels/fillers embeddings from BERT model (MASK) under one context setting.
@@
Table 6: Maximum silhouette scores on the k-means clustering of the top 15 selected English backchannels/fillers when using the LLaMA-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Discourse Marker no_ft_no_ctx
ft_no_ctx
(next-token)
no_ft_one_ctx
ft_one_ctx
(next-token)
no_ft_full_ctx
ft_full_ctx
(next-token)
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Discourse Marker no_ft_no_ctx ft_no_ctx no_ft_one_ctx ft_one_ctx no_ft_full_ctx ft_full_ctx
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Discourse Marker no_ft_no_ctx
ft_no_ctx
(next-token)
ft_no_ctx
(turn-taking)
no_ft_one_ctx
ft_one_ctx
(next-token)
ft_one_ctx
(turn-taking)
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Discourse Marker no_ft_no_ctx
ft_no_ctx
no_ft_one_ctx
ft_one_ctx
ori 100 ori 100 ori 100 ori 100
uh 0.200 0.214 0.086 0.086 0.188 0.167 0.071 0.090
yeah 0.145 0.175 0.147 0.181 0.166 0.156 0.101 0.168
uh-huh 0.148 0.195 0.172 0.197 0.173 0.184 0.115 0.268
well 0.129 0.187 0.167 0.217 0.135 0.159 0.205 0.243
right 0.104 0.132 0.162 0.211 0.140 0.140 0.256 0.305
oh 0.121 0.151 0.171 0.209 0.154 0.151 0.134 0.166
um 0.199 0.221 0.164 0.143 0.195 0.179 0.082 0.111
okay 0.120 0.138 0.150 0.190 0.123 0.124 0.163 0.202
no 0.122 0.169 0.169 0.218 0.128 0.120 0.190 0.240
yes 0.138 0.171 0.140 0.163 0.149 0.140 0.156 0.191
so 0.149 0.201 0.209 0.267 0.141 0.155 0.181 0.280
oh yeah 0.159 0.192 0.157 0.189 0.179 0.164 0.108 0.133
huh 0.214 0.286 0.209 0.246 0.200 0.180 0.098 0.128
mmhmm 0.178 0.239 0.187 0.251 0.137 0.196 0.097 0.128
of course 0.154 0.196 0.093 0.122 0.149 0.158 0.078 0.110
Average ↑ 0.152 0.191 0.159 0.192 0.157 0.158 0.135 0.185
@@
Discourse Marker no_ft_no_ctx
ft_no_ctx
(next-token)
no_ft_one_ctx
ft_one_ctx
(next-token)
no_ft_full_ctx
ft_full_ctx
(next-token)
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Discourse Marker no_ft_no_ctx ft_no_ctx no_ft_one_ctx ft_one_ctx no_ft_full_ctx ft_full_ctx
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Discourse Marker no_ft_no_ctx
ft_no_ctx
(next-token)
ft_no_ctx
(turn-taking)
no_ft_one_ctx
ft_one_ctx
(next-token)
ft_one_ctx
(turn-taking)
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Discourse Marker no_ft_no_ctx
ft_no_ctx
no_ft_one_ctx
ft_one_ctx
ori 100 ori 100 ori 100 ori 100
うん (un) 0.204 0.229 0.377 0.436 0.231 0.211 0.358 0.385
あ (a) 0.134 0.182 0.405 0.443 0.159 0.149 0.370 0.397
はい (hai) 0.241 0.176 0.194 0.376 0.151 0.141 0.159 0.456
え (e) 0.166 0.242 0.302 0.428 0.131 0.244 0.348 0.420
そう (sou) 0.207 0.184 0.362 0.427 0.221 0.140 0.353 0.420
ま (ma) 0.140 0.138 0.385 0.457 0.129 0.070 0.365 0.426
なんか (nanka) 0.094 0.118 0.367 0.103 0.078 0.102 0.357 0.120
あの (ano) 0.113 0.116 0.382 0.299 0.122 0.092 0.373 0.289
ん (n) 0.153 0.107 0.329 0.237 0.191 0.082 0.349 0.263
そうです (soudesu) 0.244 0.136 0.318 0.345 0.133 0.119 0.348 0.159
は (ha) 0.186 0.135 0.393 0.274 0.208 0.121 0.336 0.292
ね (ne) 0.220 0.091 0.327 0.284 0.197 0.074 0.321 0.288
いや (iya) 0.152 0.154 0.390 0.178 0.164 0.123 0.377 0.172
へー (he-) 0.168 0.187 0.139 0.341 0.100 0.104 0.082 0.258
そうか (souka) 0.246 0.158 0.343 0.254 0.290 0.134 0.355 0.204
Average ↑ 0.178 0.208 0.334 0.401 0.167 0.180 0.323 0.383
(a) (a) no_ft_one_context
(b) (b) ft_one_context (next_token)
Figure 17: Distance matrices of top 15 English backchannels/fillers in LLaMA-3 model before and after fine-tuning.
(a) (a) no_ft_one_context
(b) (b) ft_one_context (next_token)
Figure 18: Distance matrices of top 15 English backchannels/fillers in Qwen-3 model before and after fine-tuning.
(a) (a) no_ft_one_context
(b) (b) ft_one_context (next_token)
Figure 19: Distance matrices of top 15 English backchannels/fillers in GPT-2 model before and after fine-tuning.
(a) (a) no_ft_one_context
(b) (b) ft_one_context (masking)
Figure 20: Distance matrices of top 15 English backchannels/fillers in BERT model before and after fine-tuning.
(a) (a) no_ft_one_context
(b) (b) ft_one_context (next_token)
Figure 21: Distance matrices of top 15 Japanese backchannels/fillers in LLaMA-3 model before and after fine-tuning.
(a) (a) no_ft_one_context
(b) (b) ft_one_context (next_token)
Figure 22: Distance matrices of top 15 Japanese backchannels/fillers in Qwen-3 model before and after fine-tuning.
(a) (a) no_ft_one_context
(b) (b) ft_one_context (next_token)
Figure 23: Distance matrices of top 15 Japanese backchannels/fillers in GPT-2 model before and after fine-tuning.
(a) (a) no_ft_one_context
(b) (b) ft_one_context (masking)
Figure 24: Distance matrices of top 15 Japanese backchannels/fillers in BERT model before and after fine-tuning.
Figure 25: k value change before and after fine-tuning from different LM’s results extracted from the clustering analysis.In general, there is a slight increasing of k after fine-tuning among the different fine-tuning strategies. Exceptions are LLaMa3 for Japanese under the no-context setting; Qwen model under the no-context setting for both language;
Table 14: K values of optimal results on the k-means clustering of the top 15 selected English backchannels/fillers when using the LLaMA-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA. no_ft_no_ctx: no fine-tuning with no context information; ft_no_ctx: fine-tuning with no context information; no_ft_one_ctx: no fine-tuning with one context; ft_one_ctx: fine-tuning with one context; no_ft_full_ctx: no fine-tuning with full context; ft_full_ctx: fine-tuning with full context.
Discourse Marker no_ft_no_ctx
ft_no_ctx
(next-token)
no_ft_one_ctx
ft_one_ctx
(next-token)
no_ft_full_ctx
ft_full_ctx
(next-token)
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Table 15: K values of optimal results on the k-means clustering of the top 15 selected English backchannels/fillers when using the Qwen-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Discourse Marker no_ft_no_ctx ft_no_ctx no_ft_one_ctx ft_one_ctx no_ft_full_ctx ft_full_ctx
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Table 16: K values of optimal results on the k-means clustering of the top 15 selected English backchannels/fillers when using the GPT-2 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Discourse Marker no_ft_no_ctx
ft_no_ctx
(next-token)
ft_no_ctx
(turn-taking)
no_ft_one_ctx
ft_one_ctx
(next-token)
ft_one_ctx
(turn-taking)
ori 100 ori 100 ori 100 ori 100 ori 100 ori 100
@@
Table 17: K values of optimal results on the k-means clustering of the top 15 selected English backchannels/fillers when using the BERT model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Discourse Marker no_ft_no_ctx
ft_no_ctx
no_ft_one_ctx
ft_one_ctx
ori 100 ori 100 ori 100 ori 100

### v2

@@ D.4 Supporting Details for NLG Evaluation Result Shown in Figure 6
Table 4: Evaluation metrics for generated backchannels/fillers in the English NLG task.
@@
Perplexity $\Downarrow$ 197.67 $\pm$ 84.75 5.30 $\pm$ 1.10 202.06 $\pm$ 92.69 91.32 $\pm$ 11.70 158.64 $\pm$ 72.19 6.98 $\pm$ 1.51
BERTScore (F1 %) $\uparrow$ 78.69% $\pm$ 0.13% 79.99% $\pm$ 0.04% 76.02% $\pm$ 0.06% 79.73% $\pm$ 0.05% 79.67% $\pm$ 0.05% 79.99% $\pm$ 0.04%
BLEUScore $\uparrow$ 0.0600 $\pm$ 0.0012 0.0697 $\pm$ 0.0012 0.0698 $\pm$ 0.0012 0.0800 $\pm$ 0.0013 0.0544 $\pm$ 0.0011 0.0731 $\pm$ 0.0012
Table 5: Evaluation metrics for generated backchannels/fillers in the Japanese NLG task.
@@
Perplexity $\Downarrow$ 255.94 $\pm$ 72.48 28.51 $\pm$ 4.58 748.63 $\pm$ 31.24 90.17 $\pm$ 22.50 195.36 $\pm$ 93.86 53.51 $\pm$ 28.74
BERTScore (F1 %) $\uparrow$ 62.68% $\pm$ 0.06% 66.31% $\pm$ 0.04% 63.55% $\pm$ 0.07% 63.39% $\pm$ 0.04% 62.96% $\pm$ 0.05% 63.75% $\pm$ 0.05%
BLEUScore $\uparrow$ 0.00005 $\pm$ 0.00005 0.00030 $\pm$ 0.00010 0.00021 $\pm$ 0.00010 0.00036 $\pm$ 0.00010 0.00019 $\pm$ 0.00010 0.00026 $\pm$ 0.00005
D.5 Additional t-SNE Visualisations
Figure 3 in Section 4 shows t-SNE plots of how the embeddings of backchannels/fillers in the LLaMA-3 model change with fine-tuning in the one-context setting. In Figures 10 to 19 we provide additional t-SNE visualisations for the other models and settings. One of the general trends which can be observed is that including more contextual information can lead to better representation of backchannels/fillers (in terms of clearer borders among different ones.). For example, based on English data from Figures 10 and 11 (no-context vs. full-context), Figures 12 and 13 (no-context vs. one-context), it can be concluded that including contextual information in fine-tuning leads to better t-SNE visualisation. However, there also exist exceptions. In the Japanese data in Figures 13 and 14 (one-context vs. full-context), for example, increasing context information for fine-tuning results in poorer t-SNE visualisations.
Another issue that we noticed is that for the BERT models the improvement of the backchannels/fillers representation is not reflected well in the t-SNE visualisations: In comparison to modern LLMs (such as LLaMA-3 and Qwen-3), BERT is much more lightweight and supposedly saw less backchannels/fillers during its pre-training. However, according to Figures 18 and 19, the border among different backchannels/fillers is already quite clear before fine-tuning. This concern is not further discussed here, and could be studied in future work.
D.6 Statistics of Silhouette Scores and K-Means Values Under PCA
To ensure the statistical robustness of our evaluation, we further estimate the average silhouette scores using a bootstrap resampling approach ($n=1000$) (Efron, 1992). As summarized in Table  6, we sample the embeddings with replacement and report the mean silhouette score alongside its 95% bootstrap confidence interval half-width. Tables 7 to 14 respectively show the silhouette scores of each backchannel/filler in both the original dimensional space and after dimensionality reduction to 100 dimensions using Principal Component Analysis (PCA), along with the average silhouette scores for English and Japanese data. A general trend is that, with fine-tuning the silhouette score increases for most of the cases. When there is no fine-tuning, adding context size (no_ctx to one_ctx to full_ctx) will also have the same effect. However, combining context size with fine-tuning as well as increasing context size will lower the silhouette score. Additionally, Figure 20 and Tables 15 to 22 summarize how the $k$-value changes with fine-tuning.
D.7 Additional Confusion Matrices
Figure 5 in Section 4 shows the confusion matrices of distances among the top 15 Japanese backchannels/fillers in the Qwen-3 model before and after fine-tuning. Here, we provide additional confusion matrices for English (LLaMA-3, Qwen-3, GPT-2, and BERT) in Figures 21 to 24 and Japanese (LLaMA-3, GPT-2, and BERT) in Figures 25 to 27.
In the confusion matrices, darker colours indicate greater distances and lighter colours indicate greater similarity. The confusion matrices for English and Japanese show that for the more recent LLM models (LLaMA-3 and Qwen-3), fine-tuning makes the differences among selected backchannels/fillers more pronounced. Change is less evident for GPT-2 and BERT.
@@ D.7 Additional Confusion Matrices
Figure 10: t-SNE plots of the backchannels/fillers embeddings from the LLaMA-3 model (NTP). Setting: no-context.
@@
Figure 11: t-SNE plots of the backchannels/fillers embeddings from the LLaMA-3 model (NTP). Setting: full-context.
@@
Figure 12: t-SNE plots of the backchannels/fillers embeddings from the Qwen-3 model (NTP). Setting: no-context.
@@
Figure 13: t-SNE plots of the backchannels/fillers embeddings from the Qwen-3 model (NTP). Setting: one-context.
@@
Figure 14: t-SNE plots of the backchannels/fillers embeddings from the Qwen-3 model (NTP). Setting: full-context.
@@
Figure 15: t-SNE plots of the backchannels/fillers embeddings from the GPT-2 model (NTP). Setting: no-context.
@@
Figure 16: t-SNE plots of the backchannels/fillers embeddings from the GPT-2 model (NTP). Setting: one-context.
@@
Figure 17: t-SNE plots of the backchannels/fillers embeddings from the GPT-2 model (TTP). Setting: no- and one-context.
@@
Figure 18: t-SNE plots of the backchannels/fillers embeddings from the BERT model (MASK). Setting no-context.
@@
Figure 19: t-SNE plots of the backchannels/fillers embeddings from the BERT model (MASK): Setting: one-context.
Table 6: Evaluation results across different models and tasks. Results are reported as mean $\pm$ 95% bootstrap confidence interval half-width. ‘Base’ denotes the original model without fine-tuning, and ‘FT’ denotes the fine-tuned model. Full-context evaluation is exclusive to LLaMA-3 and Qwen-3.
Model Task Lang No-Context One-Context Full-Context
Base FT Base FT Base FT
BERT MASK EN 0.144 $\pm$ 0.011 0.241 $\pm$ 0.008 0.213 $\pm$ 0.008 0.391 $\pm$ 0.017
JP 0.213 $\pm$ 0.008 0.391 $\pm$ 0.017 0.197 $\pm$ 0.004 0.429 $\pm$ 0.012
GPT-2 NTP EN 0.274 $\pm$ 0.012 0.328 $\pm$ 0.021 0.149 $\pm$ 0.002 0.311 $\pm$ 0.020
JP 0.157 $\pm$ 0.004 0.288 $\pm$ 0.010 0.101 $\pm$ 0.003 0.273 $\pm$ 0.008
TTP EN 0.289 $\pm$ 0.019 0.211 $\pm$ 0.018
JP 0.284 $\pm$ 0.014 0.261 $\pm$ 0.014
LLaMA-3 NTP EN 0.450 $\pm$ 0.009 0.588 $\pm$ 0.011 0.183 $\pm$ 0.002 0.291 $\pm$ 0.027 0.210 $\pm$ 0.020 0.301 $\pm$ 0.018
JP 0.257 $\pm$ 0.009 0.450 $\pm$ 0.014 0.179 $\pm$ 0.003 0.335 $\pm$ 0.010 0.318 $\pm$ 0.021 0.408 $\pm$ 0.037
Qwen-3 NTP EN 0.253 $\pm$ 0.013 0.379 $\pm$ 0.065 0.157 $\pm$ 0.003 0.292 $\pm$ 0.021 0.189 $\pm$ 0.020 0.322 $\pm$ 0.028
JP 0.172 $\pm$ 0.011 0.452 $\pm$ 0.068 0.154 $\pm$ 0.002 0.263 $\pm$ 0.014 0.173 $\pm$ 0.004 0.181 $\pm$ 0.012
@@
Table 7: Maximum silhouette scores on the $k$-means clustering of the top 15 selected English backchannels/fillers when using the LLaMA-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA (no_ft: no fine-tuning; ft: fine-tuning; Settings: no_ctx: no-context; one_ctx: one-context; full_ctx: full-context)
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) no_ft, one_ctx ft, one_ctx (NTP) no_ftm full_ctx ft, full_ctx (NTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) no_ft, one_ctx ft, one_ctx (NTP) no_ft, full_ctx ft, full_ctx (NTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) ft, no_ctx (TTP) no_ft, one_ctx ft, one_ctx (NTP) ft, one_ctx (TTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Backchannel/Filler no_ft, no_ctx ft, no_ctx (MASK) no_ft, one_ctx ft, one_ctx (MASK)
orig. 100 orig. 100 orig. 100 orig. 100
uh 0.097 0.122 0.083 0.103 0.172 0.189 0.054 0.074
yeah 0.085 0.118 0.216 0.245 0.098 0.119 0.079 0.104
uh-huh 0.104 0.129 0.267 0.343 0.121 0.148 0.119 0.156
well 0.072 0.095 0.211 0.252 0.050 0.064 0.186 0.233
right 0.095 0.127 0.251 0.313 0.110 0.153 0.354 0.426
oh 0.079 0.097 0.225 0.256 0.090 0.113 0.075 0.093
um 0.091 0.117 0.144 0.178 0.192 0.206 0.078 0.112
okay 0.088 0.114 0.219 0.216 0.074 0.097 0.175 0.218
no 0.093 0.117 0.190 0.242 0.067 0.086 0.195 0.255
yes 0.096 0.122 0.206 0.227 0.071 0.089 0.122 0.146
so 0.144 0.180 0.177 0.220 0.069 0.093 0.219 0.291
oh yeah 0.115 0.142 0.162 0.178 0.119 0.139 0.114 0.137
huh 0.229 0.294 0.242 0.283 0.103 0.126 0.189 0.194
mmhmm 0.205 0.241 0.270 0.331 0.103 0.109 0.093 0.132
of course 0.069 0.096 0.088 0.111 0.050 0.073 0.088 0.103
Average ↑ 0.111 0.141 0.197 0.233 0.099 0.120 0.143 0.178
@@
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) no_ft, one_ctx ft, one_ctx (NTP) no_ft, full_ctx ft, full_ctx (NTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) no_ft, one_ctx ft, one_ctx (NTP) no_ft, full_ctx ft, full_ctx (NTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) ft, no_ctx (TTP) no_ft, one_ctx ft, one_ctx (NTP) ft, one_ctx (TTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Backchannel/Filler no_ft, no_ctx ft, no_ctx (MASK) no_ft, one_ctx ft, one_ctx (MASK)
orig. 100 orig. 100 orig. 100 orig. 100
うん (un) 0.232 0.265 0.385 0.459 0.211 0.242 0.463 0.487
あ (a) 0.147 0.166 0.399 0.393 0.141 0.171 0.447 0.428
はい (hai) 0.288 0.330 0.423 0.195 0.209 0.240 0.440 0.213
え (e) 0.164 0.191 0.293 0.364 0.150 0.170 0.406 0.425
そう (sou) 0.217 0.258 0.379 0.419 0.205 0.226 0.403 0.495
ま (ma) 0.148 0.177 0.400 0.430 0.136 0.160 0.414 0.491
なんか (nanka) 0.105 0.134 0.329 0.372 0.115 0.143 0.406 0.455
あの (ano) 0.117 0.157 0.342 0.386 0.114 0.145 0.372 0.425
ん (n) 0.155 0.192 0.429 0.488 0.161 0.190 0.523 0.569
そうです (soudesu) 0.136 0.170 0.227 0.303 0.119 0.140 0.351 0.393
は (ha) 0.208 0.250 0.403 0.437 0.204 0.247 0.431 0.484
ね (ne) 0.206 0.264 0.324 0.371 0.169 0.210 0.377 0.424
いや (iya) 0.162 0.185 0.408 0.461 0.165 0.185 0.462 0.511
へー (he-) 0.160 0.176 0.249 0.287 0.152 0.180 0.081 0.104
そうか (souka) 0.203 0.237 0.346 0.375 0.258 0.296 0.410 0.436
Average ↑ 0.177 0.210 0.356 0.383 0.167 0.196 0.399 0.423
Figure 20: Change in the $k$-value before and after fine-tuning, as extracted from the clustering analysis of different LMs’ results. A slight increasing of $k$ after fine-tuning can generally be observed among the different fine-tuning strategies. Exceptions are the no-context setting for the LLaMA-3 model for Japanese and the Qwen-3 model for both languages. It is worth noting that a decrease of the $k$-value does not necessarily indicate failure, but can also mean that the clustering effect is more salient after fine-tuning.
Table 15: $k$-values of optimal results on the $k$-means clustering of the top 15 selected English backchannels/fillers when using the LLaMA-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) no_ft, one_ctx ft, one_ctx (NTP) no_ft, full_ctx ft, full_ctx (NTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Table 16: $k$-values of optimal results on the $k$-means clustering of the top 15 selected English backchannels/fillers when using the Qwen-3 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) no_ft, one_ctx ft, one_ctx (NTP) no_ft, full_ctx ft, full_ctx (NTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Table 17: $k$-values of optimal results on the $k$-means clustering of the top 15 selected English backchannels/fillers when using the GPT-2 model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Backchannel/Filler no_ft, no_ctx ft, no_ctx (NTP) ft, no_ctx (TTP) no_ft, one_ctx ft, one_ctx (NTP) ft, one_ctx (TTP)
orig. 100 orig. 100 orig. 100 orig. 100 orig. 100 orig. 100
@@
Table 18: $k$-values of optimal results on the $k$-means clustering of the top 15 selected English backchannels/fillers when using the BERT model in both the original dimensional space and after dimensionality reduction to 100 dimensions using PCA.
Backchannel/Filler no_ft, no_ctx ft, no_ctx (MASK) no_ft, one_ctx ft, one_ctx (MASK)
orig. 100 orig. 100 orig. 100 orig. 100